# Beyond Check-in Counts: Redefining Popularity for POI Recommendation with Users and Recency

### v1

@@ Abstract.
The next POI (point of interest) recommendation aims to predict users’ immediate future movements based on their prior records and present circumstances, which will be very beneficial to service providers as well as users. The popularity of the POI over time is one of the primary deciding factors for choosing the next POI to visit. The majority of research in recent times has paid more attention to the number of check-ins to define the popularity of a point of interest, disregarding the temporal impact or number of people checking in for a particular POI. In this paper, we propose a recency-oriented definition of popularity that takes into account the temporal effect on POI’s popularity, the number of check-ins, as well as the number of people who registered those check-ins. Thus, recent check-ins get prioritized with more weight compared to the older ones. Experimental results demonstrate that performance is better with recency-aware popularity definitions for POIs than with solely check-in count-based popularity definitions.
POI recommendation, POI Popularity, Recency, Temporal Check-in
copyright: acmcopyrightjournalyear: 2018doi: 10.1145/1122445.1122456conference: Woodstock ’18: ACM Symposium on Neural Gaze Detection; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYprice: 15.00isbn: 978-1-4503-$9$-9/18/06ccs: Information systems Recommender systemsccs: Information systems Collaborative filteringccs: Information systems Recommendation algorithms
1. Introduction
@@
Over the past few years, location-based social networks (LBSNs) have experienced significant developments, including Yelp, Foursqu-are, and others. By checking-in at places of interests, users may share their locations and experiences with their friends. The visited point of interest (POI) together with any accompanying contexts (e.g., categories, timestamp, GPS) that define the user’s movement are typically included in a check-in record. The enormous amount of check-in data produced by millions of users in LBSNs offers a fantastic chance to investigate the underlying patterns in user check-in behavior. Numerous point-of-interest and travel recommendation systems take advantage of this vast quantity of information.
@@
When choosing the next point of interest (POI), nearly all trip and POI recommender systems employ popularity as one of their measures. The majority of them characterize the number of check-ins as a POI’s popularity indicator. However, determining popularity only by check-in numbers isn’t necessarily the best course of action. This provides us with a few things to improve. First, a point’s level of popularity varies with time. Points A and B, for instance, have nearly equal numbers of check-ins. While the majority of the check-ins in point-B were done a few years ago, the majority of the check-ins in point-A were made recently. Point-B was apparently popular back then, based on this scenario, but choosing point-A over point-B makes a lot more sense now. Second, a place’s popularity among a certain demographic does not guarantee that it will be favored by everyone. For instance, a significant number of individuals go to point-A regularly, let’s say on a weekly basis, whereas only a handful of people meet at place-B but they go there every day. Hence, even though there are exactly the same number of check-ins at these two locations, it is very possible that a newcomer would select point-A over point-B because point-A is preferred by a wider range of individuals. Our next POI selection problem will include redefining popularity in light of these observations. Our suggested strategy essentially incorporates the new definition of popularity into the GETNext method (Yang et al., 2022) which builds a graph enhanced transformer model by combining the user’s general preference, spatio-temporal context, global transition pattern, and time-aware category embedding to predict the user’s next move. To extract the relevant information from the raw data, we first preprocessed the data. Next, using the user’s current trajectory as a guide, we apply the modified version of the GETNext algorithm to the processed data to determine the next top-k POI recommendation. In short, the following sums up our contributions:
@@
2. Related Work
@@
For POI or trip recommendation systems, there are numerous approaches available, ranging from matrix factorization to recurrent neural networks (RNN) to graph-based techniques. Markov chains, which have been widely used in other sequential recommendation problems, are among the strategies used in early research. For instance, by using a matrix factorization technique that combines FPMC (Personalized Markov Chains) (Rendle et al., 2010), Cheng et al. (Cheng et al., 2013) proposed next points of interest in their pioneering study. Similar to this, Zhang et al. (Zhang et al., 2014) suggested modeling the sequential transitive influence using an additive Markov chain. These early methods fall short of deep neural network models when it comes to modeling sequence data.
@@
Recently, research has focused on deep learning and advanced embedding methods (Feng et al., 2020). To represent the temporal dynamics and sequential correlations, RNN variants were proposed (Liu et al., 2016; Sun et al., 2020; Wu et al., 2019, 2020; Zhao et al., 2020, 2019). Liu et al. (2016) presented spatial temporal recurrent neural networks (ST-RNN) (Liu et al., 2016). Spatial and temporal contexts were incorporated into RNN layers by these networks. In particular, geographic distance transition matrices characterize spatial contexts, whereas time transition matrices reflect temporal context. LSTM was also utilized to approximate the short- and long-term preferences of the users. In order to acquire user preferences and a standard LSTM model for short-term trajectory mining, the creators of PLSPL (Wu et al., 2020) and LSPL (Wu et al., 2019) trained a general embedding layer. For the final recommendation, DeepMove (Feng et al., 2018) suggested using a recurrent neural network to simulate the sequential transitions and an attention model to capture the multi-level periodicity pattern.
@@
A powerful paradigm is provided by graph-based methods, like those that make use of location-based social networks (LBSN), especially for conventional (non-sequential) recommendation tasks. A tripartite graph including POI, session, and user nodes is the geographical-temporal influences aware graph (GTAG), constructed by Yuan et al (Yuan et al., 2014). To train an embedding of the current POI, the authors of a recent work (Li et al., 2021) randomly sampled previous and next check-ins from different check-in sequences for each POI in the given check-in sequence. The embedding does not specifically reflect the global (multi-hop) transition patterns because the goal was to capture local (one-hop) transition of POIs. Graph-based techniques were employed by Yang et al. (Yang et al., 2022) in their next POI recommendation. The authors specifically used a uniform graph structure in their strategy to recommend the next POI. Global patterns among all POIs were manifested in a unified graph structure, which was their modeled trajectory flow map. In the next POI recommendation problem, the authors employed graph-based learning to encode generic transitional information about POIs.
It is our finding that a POI’s popularity is greatly influenced by the number of users checking in and the recentness of check-in records. The majority of the recent research methods solely took into account the quantity of check-ins, giving less thought to the number of users checking in. We offer a new definition of popularity for a POI that takes into account the total number of users who check in as well as the number of check-ins. Additionally, it gives precedence to the current trend in popularity above those that were successful in the past but aren’t as popular currently.
3. Problem Formulation
@@
Now, let’s explore the key ideas presented in the paper. Starting with the concept of users, represented by $U$, we create a collection $U=\{u_{1},u_{2},\ldots,u_{M}\}$, in which $M$ is the total number of users. Points of interest (POIs) are the next category, and they include different locations such as restaurants or hotels or coffee shops or parks or malls or clothing stores or bus stations or airports etc. They are represented by the symbol $P=\{p_{1},p_{2},\ldots,p_{N}\}$, where $N$ denotes the number of unique POIs. The timestamps, which represent discrete points in time, are stored in the set $T=\{t_{1},t_{2},\ldots,t_{K}\}$, where $K$ represents the total timestamps.
[…]
@@
The sequence of a user’s check-in activities, $Q_{u}=(q_{u}^{1},q_{u}^{2},q_{u}^{3},\ldots)$, is an ordered collection of these events, and each $q_{u}^{i}$ represents the $i$-th check-in record.
Definition 3.2 (Check-in Set): The check-in sequences for each user are included in the collection/set is denoted as $Q_{U}=\{Q_{u_{1}},Q_{u_{2}},\ldots,Q_{u_{M}}\}$.
We divide each user’s check-in sequence $Q_{u}$ into successive trajectories during data preprocessing, which are represented as $Q_{u}=S_{u}^{1}\oplus S_{u}^{2}\oplus\ldots$, concatenating them with $\oplus$. These trajectories, which differ in length, represent a sequence of check-ins that take place over predetermined periods of time, like an entire day.
@@
4. Revised GETNext Method for the Next POI Recommendation
We modified the GETNext model (Yang et al., 2022) by introducing a new measurement for POI popularity. In our case, popularity is represented by a multivariate formula that attempts to capture different aspects of user interaction. Now let’s explore its specifics:
(1) Popularity = $\begin{split}\text{Popularity = }&\beta\left(\alpha\cdot C_{\text{user}}^{\text{recent}}+(1-\alpha)\cdot C_{\text{check-in}}^{\text{recent}}\right)\\ &+(1-\beta)\left(\alpha\cdot C_{\text{user}}^{\text{past}}+(1-\alpha)\cdot C_{\text{check-in}}^{\text{past}}\right)\end{split}$
@@ 4. Revised GETNext Method for the Next POI Recommendation
Figure 1. An overview of the GETNext model. (Yang et al., 2022)
4.1. Learning with Trajectory Flow Map
4.1.1. Understanding Trajectory Flow Map
@@ 4.1.2. Learning POI Embedding
4.1.3. Deciphering Transition Attention Map
@@
4.2. Contextual Embedding Module
@@
4.2.1. POI-User Embeddings Fusion
@@
(2) $e_{u}=\text{fembed}(u)\in\mathbb{R}^{\Omega}.$
@@
(3) $e_{p,u}=\sigma(w_{p,u}[e_{p};e_{u}]+b_{p,u})\in\mathbb{R}^{\Omega\times 2},$
@@
4.2.2. Time-Category Embeddings Fusion
@@
(4) $e_{c,t}=\sigma(w_{c,t}[e_{t};e_{c}]+b_{c,t})\in\mathbb{R}^{\Psi\times 2},$
@@
4.3. Transformer Encoder and MLP Decoders
4.3.1. Transformer Encoder
@@
(7) $\text{head}_{1}=S^{\prime}X^{[l]}W_{v}\in\mathbb{R}^{k\times\frac{d}{h}}$
@@
(8) $Multihead(X^{[l]})=[head_{1};\ldots;head_{h}]\times W_{o}\in\mathbb{R}^{k\times d}$
@@
(9) $X^{[l]}_{\text{attn}}=\text{LayerNorm}(X^{[l]}+\text{Multihead}(X^{[l]}))$
@@
(10) $X^{[l]}_{\text{FC}}=\text{ReLU}(W_{1}X^{[l]}_{\text{attn}}+b_{1})W_{2}+b_{2}\in\mathbb{R}^{k\times d}$
@@
(11) $X^{[l+1]}=\text{LayerNorm}(X^{[l]}_{\text{attn}}+X^{[l]}_{\text{FC}})\in\mathbb{R}^{k\times d}$
@@ 4.3.2. MLP Decoders
(12) $\hat{Y}_{\text{poi}}=X^{[l^{*}]}W_{\text{poi}}+b_{\text{poi}}$
@@
(13) $\hat{Y}_{\text{time}}=X^{[l^{*}]}W_{\text{time}}+b_{\text{time}}$
@@
(14) $\hat{Y}_{\text{cat}}=X^{[l^{*}]}W_{\text{cat}}+b_{\text{cat}}$
@@ 4.3.3. Loss
(15) $L_{\text{final}}=L_{\text{poi}}+10\times L_{\text{time}}+L_{\text{cat}}$
@@
5. Experiment
5.1. Dataset
As part of our experiment, we carefully examined the FourSquare-NYC public dataset, which is an extensive collection of urban check-in patterns in New York City that was carefully collected and described by Dingqi et al. (Yang et al., 2014). Covering the busy streets of April 2012 to the quiet streets of February 2013, this dataset captures a variety of user interactions within the dynamic cityscape. Every entry in this list represents an assembly of essential data, such as the user’s name, the point of interest (POI) that was visited, the particular category that the POI falls under, the exact GPS coordinates indicating its location, and the timestamp indicating the interaction.
We started the process of data refining in order to obtain more sophisticated insights through eliminating users who had very few check-ins in the past, leaving only those who had an important record of at least ten actions that were recorded. In a similar manner, we removed POIs with less than 10 check-in records from the dataset to make sure our analysis was supported by reliable and significant data points.
@@ 5.1. Dataset
Table 1. Dataset Statistics
@@
Table 2. Experimental Results
$\alpha$ $\beta$ Acc@1 Acc@5 Acc@10 Acc@20 MRR

### v2

@@ Abstract
The task of point-of-interest (POI) recommendation is to predict users’ immediate future movements based on their previous records and present circumstances. Popularity is considered as one of the primary deciding factors for selecting the next place to visit. Existing approaches mainly focused on the number of check-ins to model the popularity of a POI. However, not enough attention is paid to the temporal impact or number of people check-ins for a particular POI. Thus, to prioritize more on recent check-ins, we propose recency-oriented definition of POI’s popularity by considering the temporal effect of the POIs, the number of check-ins, as well as the number of users who registered in those check-ins. Our experimental results on real dataset show the efficacy of the proposed approach.
Keywords:
POI recommendation POI Popularity Recency Temporal Effect.
1 Introduction
@@
Over the past few years, location-based social networks (LBSNs) have experienced significant developments, including Yelp, Foursquare, and others. By checking-in at places of interests, users may share their locations and experiences with their friends. The visited point of interest (POI) together with any accompanying contexts (e.g., categories, timestamp, GPS) that define the user’s movement are typically included in a check-in record. The enormous amount of check-in data produced by millions of users in LBSNs offers a fantastic chance to investigate the underlying patterns in user check-in behavior. Numerous point-of-interest and travel recommendation systems take advantage of this vast quantity of information.
When choosing the next point of interest (POI), nearly all trip and POI recommender systems employ popularity as one of their measures. The majority of them characterize the number of check-ins as a POI’s popularity indicator. However, determining popularity only by check-in numbers isn’t necessarily the best course of action. This provides us with a few things to improve.
First, a point’s level of popularity varies with time. Points A and B, for instance, have nearly equal numbers of check-ins. While the majority of the check-ins in point-B were done a few years ago, the majority of the check-ins in point-A were made recently. Point-B was apparently popular back then, based on this scenario, but choosing point-A over point-B makes a lot more sense now.
Second, a place’s popularity among a certain demographic does not guarantee that it will be favored by everyone. For instance, a significant number of individuals go to point-A regularly, let’s say on a weekly basis, whereas only a handful of people meet at place-B but they go there every day. Hence, even though there are exactly the same number of check-ins at these two locations, it is very possible that a newcomer would select point-A over point-B because point-A is preferred by a wider range of individuals.
@@
Our next POI selection problem will include redefining popularity in light of these observations. Our suggested strategy essentially incorporates the new definition of popularity into the GETNext method [11] which builds a graph enhanced transformer model by combining the user’s general preference, spatio-temporal context, global transition pattern, and time-aware category embedding to predict the user’s next move. To extract the relevant information from the raw data, we first preprocessed the data. Next, using the user’s current trajectory as a guide, we apply the modified version of the GETNext algorithm to the processed data to determine the next top-k POI recommendation. In short, the following sums up our contributions:
@@
In Section 2, we reviewed several studies that are relevant to our work. We made an effort to summarize their approach and set it apart from ours. In-depth analysis of the subtleties of problem formulation is provided in Section 3. All of the details of our suggested methodology and initial concepts are covered in Section 4. Discussion about our experimental setup, dataset, and findings are revealed in Section 5. A summary of our overall contributions are included in Section 6 along with wrapping up the work and outlining our next research goals.
2 Related Work
@@
There are several methods available for POI or trip recommendation systems, including matrix factorization, recurrent neural networks (RNN), and graph-based approaches. Among the tactics employed in early research are Markov chains, which have been frequently used in other sequential recommendation problems. For example, Cheng et al. [1] proposed next points of interest in their groundbreaking study by combining FPMC (Personalized Markov Chains) [6] with a matrix factorization technique. In a similar vein, Zhang et al. [13] proposed to use an additive Markov chain to simulate the sequential transitive influence. When it comes to modeling sequence data, these early techniques are not as effective as deep neural network models.
@@
Deep learning and sophisticated embedding techniques have been the subject of recent study [3]. RNN variations were proposed [5, 7, 8, 9, 14, 15] to reflect the temporal dynamics and sequential correlations. Spatial temporal recurrent neural networks (ST-RNN) were introduced by Liu et al. (2016) [5]. These networks integrated temporal and spatial contexts into RNN layers. Specifically, time transition matrices describe temporal context, while geographic distance transition matrices characterize spatial contexts. The participants’ short- and long-term preferences were also approximated using LSTM. For short-term trajectory mining, the authors of PLSPL [9] and LSPL [8] trained a general embedding layer to get user preferences and a conventional LSTM model. DeepMove [2] proposed an attention model to capture the multi-level periodicity pattern and a recurrent neural network to simulate the sequential transitions.
@@
Graph-based approaches, such as those utilizing location-based social networks (LBSN), offer a potent paradigm, particularly for traditional (non-sequential) recommendation tasks. Yuan et al. [12] created the geographical-temporal influences aware graph (GTAG), a tripartite graph including POI, session, and user nodes. The authors of a recent work [4] randomly picked previous and next check-ins from different check-in sequences for each POI in the provided check-in sequence in order to train an embedding of the current POI. Since the intention was to capture local (one-hop) transition of POIs, the embedding does not particularly reflect the global (multi-hop) transition patterns. In their subsequent POI recommendation, Yang et al. [11] used graph-based approaches. To recommend the next point of interest, the authors deliberately employed a uniform graph structure in their methodology. All of the POIs’ global patterns were represented by a single graph structure, or modeled trajectory flow map. The authors used graph-based learning to encode generic transitional information about POIs in the next POI recommendation problem.
We find that the quantity of users checking in and the recentness of check-in records have a significant impact on a POI’s popularity. Most current study approaches considered only the number of check-ins and paid little attention to the number of users checking in. We propose a novel definition of popularity for a POI that considers both the total number of check-ins and the number of people who do not check-in. Furthermore, it prioritizes the current trend in popularity over those that were popular a while ago but aren’t as popular right now.
3 Problem Formulation
@@
Now, let’s explore the key ideas presented in the paper. Starting with the concept of users, represented by $U$, we create a collection $U=\{u_{1},u_{2},\ldots,u_{M}\}$, in which $M$ is the total number of users. Points of Interest (POIs) are the next category and include various locations such as restaurants, hotels, coffee shops, parks, malls, clothing stores, bus stations, airports, etc. They are represented by the symbol $P=\{p_{1},p_{2},\ldots,p_{N}\}$, where $N$ denotes the number of unique POIs. The timestamps, representing discrete points in time, are stored in the set $T=\{t_{1},t_{2},\ldots,t_{K}\}$, where $K$ represents the total number of timestamps.
Within set $P$, every POI $p$ is specified by a tuple $p=\langle\text{cat},\text{freq},\text{lat},\text{lon}\rangle$, containing its category (cat), frequency of visits (freq), and geographical coordinates (latitude (lat) and longitude (lon)). The category, denoted by cat, correlates to preset categories such as "train station" or "restaurant."
@@
The sequence of a user’s check-in activities, $Q_{u}=(q_{u}^{1},q_{u}^{2},q_{u}^{3},\ldots)$, is an ordered collection of these events, where each $q_{u}^{i}$ represents the $i$-th check-in record.
Definition 3.2 (Check-in Set): The check-in sequences for each user are included in the collection/set denoted as $Q_{U}=\{Q_{u_{1}},Q_{u_{2}},\ldots,Q_{u_{M}}\}$.
During data preprocessing, we divide each user’s check-in sequence $Q_{u}$ into successive trajectories, represented as $Q_{u}=S_{u}^{1}\oplus S_{u}^{2}\oplus\ldots$, concatenating them with $\oplus$. These trajectories, which differ in length, represent a sequence of check-ins that occur over predetermined periods, such as an entire day.
@@
4 Revised GETNext Method for the Next POI Recommendation
We modified the GETNext model [11] by introducing a new measurement for POI popularity. In our case, popularity is represented by a multivariate formula that attempts to capture different aspects of user interaction. Now let’s explore its specifics:
Popularity = $\begin{split}\text{Popularity = }&\beta\left(\alpha\cdot C_{\text{user}}^{\text{recent}}+(1-\alpha)\cdot C_{\text{check-in}}^{\text{recent}}\right)\\ &+(1-\beta)\left(\alpha\cdot C_{\text{user}}^{\text{past}}+(1-\alpha)\cdot C_{\text{check-in}}^{\text{past}}\right)\end{split}$ (1)
@@ 4 Revised GETNext Method for the Next POI Recommendation
Figure 1: An overview of the GETNext model. [11]
4.1 Learning with Trajectory Flow Map
4.1.1 Understanding Trajectory Flow Map
@@ 4.1.2 Learning POI Embedding
4.1.3 Deciphering Transition Attention Map
@@
4.2 Contextual Embedding Module
@@
4.2.1 POI-User Embeddings Fusion
@@
$e_{u}=\text{fembed}(u)\in\mathbb{R}^{\Omega}.$ (2)
@@
$e_{p,u}=\sigma(w_{p,u}[e_{p};e_{u}]+b_{p,u})\in\mathbb{R}^{\Omega\times 2},$ (3)
@@
4.2.2 Time-Category Embeddings Fusion
@@
$e_{c,t}=\sigma(w_{c,t}[e_{t};e_{c}]+b_{c,t})\in\mathbb{R}^{\Psi\times 2},$ (4)
@@
4.3 Transformer Encoder and MLP Decoders
4.3.1 Transformer Encoder
@@
$\text{head}_{1}=S^{\prime}X^{[l]}W_{v}\in\mathbb{R}^{k\times\frac{d}{h}}$ (7)
@@
$Multihead(X^{[l]})=[head_{1};\ldots;head_{h}]\times W_{o}\in\mathbb{R}^{k\times d}$ (8)
@@
$X^{[l]}_{\text{attn}}=\text{LayerNorm}(X^{[l]}+\text{Multihead}(X^{[l]}))$ (9)
@@
$X^{[l]}_{\text{FC}}=\text{ReLU}(W_{1}X^{[l]}_{\text{attn}}+b_{1})W_{2}+b_{2}\in\mathbb{R}^{k\times d}$ (10)
@@
$X^{[l+1]}=\text{LayerNorm}(X^{[l]}_{\text{attn}}+X^{[l]}_{\text{FC}})\in\mathbb{R}^{k\times d}$ (11)
@@ 4.3.2 MLP Decoders
$\hat{Y}_{\text{poi}}=X^{[l^{*}]}W_{\text{poi}}+b_{\text{poi}}$ (12)
@@
$\hat{Y}_{\text{time}}=X^{[l^{*}]}W_{\text{time}}+b_{\text{time}}$ (13)
@@
$\hat{Y}_{\text{cat}}=X^{[l^{*}]}W_{\text{cat}}+b_{\text{cat}}$ (14)
@@ 4.3.3 Loss
$L_{\text{final}}=L_{\text{poi}}+10\times L_{\text{time}}+L_{\text{cat}}$ (15)
@@
5 Experiment
5.1 Dataset
As part of our experiment, we carefully examined the FourSquare-NYC public dataset, which is an extensive collection of check-ins in New York City from April 2012 to February 2013 that was carefully collected and described by Dingqi et al. [10]. Every entry in this list represents the user’s name, the point of interest (POI) that was visited, the particular category that the POI falls under, the exact GPS coordinates indicating its location, and the timestamp indicating the interaction.
We started the process of data refining through eliminating users who had very few check-ins in the past, leaving only those who had an important record of at least ten actions that were recorded. In a similar manner, we removed POIs with less than 10 check-in records from the dataset to make sure our analysis was supported by reliable and significant data points.
@@ 5.1 Dataset
Table 1: Dataset Statistics
@@
5.2 Evaluation Metrics
[…]
$\text{MRR}=\frac{1}{m}\sum_{i=1}^{m}\frac{1}{\text{rank}}$
5.3 Results
We have conducted a number of experiments that have contributed significantly to our understanding of the complex dynamics of our model. Despite the trials, the fundamental pillars of our study, denoted by the variables $\alpha$ and $\beta$, as stated in equation 1, held firm. Our technique stayed largely unchanged across a resolute span of 20 epochs, which allowed our comparison study to remain coherent.
Check-ins sought solace in two discrete temporal domains—the recent and the past—amid the complex pattern of temporal dynamics. Those who were recorded in the history of the previous three months were labeled as recent. On the other hand, those who came before this point in time were relegated to the past.
Table 2: Experimental Results
Acc@1 Acc@5 Acc@10 Acc@20 MRR
Baseline